40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 30 00:58:19 2011, Koji, Summary, LSC, LSC whitening filter test chans_24_31_WeirdPhase.pdfOctopus.jpgTest_Inputs_Plugged_In.jpgContec_Tester_Board.jpg
    Reply  Thu Jul 7 02:23:59 2011, Jenne, Summary, LSC, LSC Whitening Filters have been fit AS55Q.png
       Reply  Thu Jul 7 15:34:44 2011, Jenne, Summary, LSC, LSC Whitening Filters have been fit 
Message ID: 4915     Entry time: Thu Jun 30 00:58:19 2011     Reply to this: 4951
Author: Koji 
Type: Summary 
Category: LSC 
Subject: LSC whitening filter test 

[Jenne, Koji]

We have tested the LSC whitening filters. In summary, they show the transfer functions mostly as expected (15Hz zerox2, 150Hz pole x2).
Only CH26 (related to the slow channel "C1:LSC-PD9_I2_WhiteGain. VAL NMS", which has PD10I label in MEDM) showed different
phase response. Could it be an anti aliasing filter bypassed???

The 32 transfer functions obtained will be fit and summarized by the ZPK parameters.


Method:

The CDS system was used in order to get the transfer functions
- For this purpose, three filter modules ("LSC-XXX_I", "LSC-XXX_Q", "LSC-XXX_DC") were added to c1lsc
in order to allow us to access to the unused ADC channels. Those filter modules have terminated outputs.
The model was built and installed. FB was restarted in order to accomodate the new channels.

- Borrow a channel from ETMY UL coil output mon. Drag the cable from the ETMY rack to the LSC analog rack.
- Use 7 BNC Ts to split the signal in to 8 SMA cables.
- Put those 8 signals into each whitening filter module.

- The excitation signal was injected to C1:SUS-ETMY_ULCOIL_EXC by AWGGUI.
- The transfer functions were measured by DTT.
- The excitation signal was filtered by the filter zpk([150;150],[15;15],1,"n")
   so that the whitened output get flat so as to ensure the S/N of the measurement.

- For the switching, we have connected the CONTEC Binary Output Test board to the BIO adapter module
   in stead of the flat cable from the BIO card. This allow us to switch the individual channels manually.

- The whitening filters of 7 channels were turned on, while the last one is left turned off.
- We believe that the transfer functions are flat and equivalent if the filters are turned off.
- Use the "off" channel as the reference and measure the transfer functions of the other channels.
- This removes the effect of the anti imaging filter at ETMY.

- Once the measurement of the 7 channels are done, switch the role of the channels and take the transfer function for the remaining one channel.

Result:

- We found the following channel assignment

  • The ADC channels and the PDs. This was known and just a confirmation. 
  • The ADC channels and the WF filter on MEDM (and name of the slow channel)

- We found that the binary IO cable at the back of the whitening filter module for ADC CH00-CH07 were not connected properly.
This was because the pins of the backplane connector were bent. We fixed the pins and the connector has been properly inserted.

- CH26 (related to the slow channel "C1:LSC-PD9_I2_WhiteGain. VAL NMS", which has PD10I label in MEDM) showed different
phase response from the others although the amplitude response is identical.

Summary of the channel assignment (THEY ARE OBSOLETE - SEPT 20, 2011)

ADC                    Whitening Filter
CH  PD                 name in medm   related slow channel name for gain
---------------------------------------------------------------------------
00  POY11I             PD1I           C1:LSC-ASPD1_I_WhiteGain. VAL NMS
01  POY11Q             PD1Q          
C1:LSC-ASPD1_Q_WhiteGain. VAL NMS
02  POX11I             PD2I           C1:LSC-SPD1_I_WhiteGain. VAL NMS
03  POX11Q             PD2Q           C1:LSC-SPD1_Q_WhiteGain. VAL NMS
04  REFL11I            PD3I           C1:LSC-POB1_I_WhiteGain. VAL NMS
05  REFL11Q            PD3Q           C1:LSC-POB1_Q_WhiteGain. VAL NMS
06  AS11I              PD4I           C1:LSC-ASPD2_I_WhiteGain. VAL NMS
07  AS11Q              PD4Q           C1:LSC-ASPD2_Q_WhiteGain. VAL NMS
08  AS55I              AS55_I         C1:LSC-ASPD1DC_WhiteGain. VAL NMS
09  AS55Q              AS55_Q         C1:LSC-SPD1DC_WhiteGain. VAL NMS
10  REFL55I            PD3_DC         C1:LSC-POB1DC_WhiteGain. VAL NMS
11  REFL55Q            PD4_DC         C1:LSC-PD4DC_WhiteGain. VAL NMS
12  POP55I             PD5_DC         C1:LSC-PD5DC_WhiteGain. VAL NMS
13  POP55Q             PD7_DC         C1:LSC-PD7DC_WhiteGain. VAL NMS
14  REFL165I           PD9_DC         C1:LSC-PD9DC_WhiteGain. VAL NMS
15  REFL165Q           PD11_DC        C1:LSC-PD11DC_WhiteGain. VAL NMS
16  NC (named XXX_I)   PD5I           C1:LSC-SPD2_I_WhiteGain. VAL NMS
17  NC (named XXX_Q)   PD5Q           C1:LSC-SPD2_Q_WhiteGain. VAL NMS
18  AS165I             PD6I           C1:LSC-SPD3_I_WhiteGain. VAL NMS
19  AS165Q             PD6Q           C1:LSC-SPD3_Q_WhiteGain. VAL NMS
20  REFL33I            PD7I           C1:LSC-POB2_I_WhiteGain. VAL NMS
21  REFL33Q            PD7Q
           C1:LSC-POB2_Q_WhiteGain. VAL NMS
22  POP22I             PD8I
           C1:LSC-ASPD3_I_WhiteGain. VAL NMS
23  POP22Q             PD8Q
           C1:LSC-ASPD3_Q_WhiteGain. VAL NMS
24  POP110I            PD9I
           C1:LSC-PD9_I1_WhiteGain. VAL NMS
25  POP110Q            PD9Q
           C1:LSC-PD9_Q1_WhiteGain. VAL NMS
26  NC (named XXX_DC)  PD10I
          C1:LSC-PD9_I2_WhiteGain. VAL NMS
27  POPDC              PD10Q
          C1:LSC-PD9_Q2_WhiteGain. VAL NMS
28  POYDC              PD11I
          C1:LSC-PD11_I_WhiteGain. VAL NMS
29  POXDC              PD11Q
          C1:LSC-PD11_Q_WhiteGain. VAL NMS
30  REFLDC             PD12I
          C1:LSC-PD12_I_WhiteGain. VAL NMS
31  ASDC               ASDC
           C1:LSC-PD12_Q_WhiteGain. VAL NMS
---------------------------------------------------------------------------

Attachment 1: chans_24_31_WeirdPhase.pdf  153 kB  Uploaded Thu Jun 30 02:34:57 2011  | Hide | Hide all
chans_24_31_WeirdPhase.pdf
Attachment 2: Octopus.jpg  102 kB  Uploaded Thu Jun 30 02:35:13 2011  | Hide | Hide all
Octopus.jpg
Attachment 3: Test_Inputs_Plugged_In.jpg  252 kB  Uploaded Thu Jun 30 02:35:26 2011  | Hide | Hide all
Test_Inputs_Plugged_In.jpg
Attachment 4: Contec_Tester_Board.jpg  55 kB  Uploaded Thu Jun 30 02:35:41 2011  | Hide | Hide all
Contec_Tester_Board.jpg
ELOG V3.1.3-